OpenStudy (anonymous):

Solve the linear system. x + y = -3 y = 2x

OpenStudy (anonymous):

@kl0723

OpenStudy (kl0723):

just plug in the y into equation one and solve for x

OpenStudy (anonymous):

y = -6?

OpenStudy (kl0723):

since y=2x, then equation one is x+ (2x) =-3 3x = -3 x = -3/3 x = -1

OpenStudy (kl0723):

now plug in the x into equation number 2... y=2x y=2(-1) ...then y=-2 :)

OpenStudy (anonymous):

oh so y = -2 and x = -2

OpenStudy (anonymous):

i mean -1

OpenStudy (kl0723):

yes :)

OpenStudy (anonymous):

medal for medal or nah

OpenStudy (anonymous):

nah^